OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 - 10. What is the value of (f + g)(-2)? -6 -3 9 -9

OpenStudy (anonymous):

plug in -2 for each function then add their results to get -9 2(-2)+1 = -3 (-2)^2 -10 = -6 -3 + -6 = -9